Company ABC has recently operated in both Spain and France. Found in 2000, ABC has grown steadily over the years to be one of the largest competitors in the ice cream business.

Exhibit 2 French Region, 2009 Results Variance Profit Plan Volume Euros (000) (000) Actual Volume Euros (000) (000) Sales

Analyze one line item in the variable costs of the French division.
Answer #1

We are going to analyse dairy ingredients (litres ) under cost of ice cream in the variable costs of the French division.

As per the the plan

For the sale of 4010 ltrs of ice cream

2887 litres of dairy ingredients are required

Therefore for 1 ltr of ice cream 0.72 litres of dairy ingredients are required.

Per unit cost estimated=7893/2887 =2.73

Actual

For the sale of 4618 ltrs of ice cream

3317 litres of dairy ingredients are required

Therefore for 1 ltr of ice cream 0.72 litres of dairy ingredients are required.

Per unit cost estimated 9142/3317 = 2.76

actual per unit cost of dairy ingredient is more than the amount planned.